# Quantifying task-relevant representational similarity using decision variable correlation

### v1

@@ Abstract
Previous studies have compared the brain and deep neural networks trained on image classification. Intriguingly, while some suggest that their representations are highly similar, others argued the opposite. Here, we propose a new approach to characterize the similarity of the decision strategies of two observers (models or brains) using decision variable correlation (DVC). DVC quantifies the correlation between decoded decisions on individual samples in a classification task and thus can capture task-relevant information rather than general representational alignment. We evaluate this method using monkey V4/IT recordings and models trained on image classification tasks. We find that model–model similarity is comparable to monkey-monkey similarity, whereas model–monkey similarity is consistently lower and, surprisingly, decreases with increasing ImageNet-1k performance. While adversarial training enhances robustness, it does not improve model–monkey similarity in task-relevant dimensions; however, it markedly increases model–model similarity. Similarly, pre-training on larger datasets does not improve model–monkey similarity. These results suggest a fundamental divergence between the task-relevant representations in monkey V4/IT and those learned by models trained on image classification tasks.
@@ 1 Introduction
Deep learning [1, 2] has revolutionized how neuroscientists construct models of the brain. For vision science, deep neural networks offer appealing candidate models for studying the primate ventral pathway [3, 4, 5] and, more recently, dorsal pathway [6, 7, 8]. A decade ago, it was reported that the internal representations in convolutional neural networks (CNNs) trained on image categorization can explain a substantial fraction of variance in higher visual areas, surpassing the classic models for these brain areas [3]. Follow-up research has tested many more variants of deep networks on their aligment with the brain using both neural data [9, 10, 11] and behavior data [12, 13, 14] from monkeys and humans. Newer deep network models are learning to leverage information with increasing efficiency and reliability, mirroring the evolutionary journey of primate brains. Naturally, one appealing hypothesis is that deep networks that exhibit higher accuracy and robustness in vision tasks, or trained on larger datasets would better explain visual processing in the brain.
The key therefore lies in how to compare deep networks and the brain.
@@
One class of methods seeks to quantify the similarity of internal representations between models and brains. This includes methods such as representational similarity analysis (RSA) analysis[15], linear regression[5, 16], and generalized shape metrics [17]. More recently, another class of methods that put more emphasize on quantifying the behavioral similarity has been proposed, including Cohen’s Kappa [13] and I2n behavioral predictivity [12, 18]. The goal of these methods is to provide a high-resolution, image-by-image comparison of the decision strategies used by neural networks and the brain. One challenge has been how to properly disentangle the model accuracy, decision biases and decision consistency from behavior [19, 13]. Interestingly, these methods seem to find contradictory trends in model-brain alignment. The reason for this remains unclear.
@@
Here, we propose a principled approach that combines the merits of model comparisons at the representation and behavior levels. Our new approach specifically quantifies the trial-by-trial consistency of two neural representations for solving a classification task, ignoring features that are irrelevant for the task. In doing so, it enables us to isolate the consistency of the decision strategy of two observers when solving a behavioral task. Our method is based on decision variable correlation (DVC) developed to measure behavioral similarity based on choice data [19], and we have generalized it to analyze the consistency of high-dimensional neural representations. Applying our method to compare neural representations for solving image recognition tasks from monkey brains and deep network models led to several interesting findings. In particular, we found that model–model and monkey–monkey similarities are comparable, whereas model–monkey similarity is consistently lower and decreases with increasing ImageNet-1k accuracy. Somewhat surprisingly, this gap is not remedied by adversarial training or training on larger datasets.
@@ 2 Background and relevant work
Community efforts have pushed towards better methods to compare brains and models and for brain-model alignment. Different factors have been hypothesized to be relevant to the alignment, including model architecture, robustness, and training data, as summarized below.
@@
Model architecture and scale One hypothesis has been that as models improve in task performance or architectural complexity, their internal representations become more brain-like. The Brain-Scores [18] of the image classification models were reported to be positively correlated to ImageNet-1k accuracy, although the trend plateaus at higher accuracy. On the other hand, studies using RSA reported that neither model scale nor architecture significantly improved alignment to human behavioral similarity judgments [9]. Another study using RSA reported negative correlation between alignment to human neural activity and model complexity (in FLOPs) [20]. Studies using Cohen’s Kappa reported that human-model behavioral consistency at the image level remains low despite improved performance on out-of-distribution datasets with scaling[13, 14].
@@
Robustness In addition to good performance, the primate visual system is robust against external and internal noise, prompting the question of whether robustness to adversarial perturbations or corruptions is related to brain-model alignment [21]. By enforcing alignment with monkey IT representations, models exhibited both enhanced adversarial robustness and increased behavioral alignment with human subjects. Another study found that model metamers – artificial stimuli that elicit the same response as natural stimuli, generated by robust models– are more recognizable to humans, but is not itself predictive of recognizability [22]. However, studies using Cohen’s Kappa report that robust models still diverge from humans on their error patterns[14].
@@
Rich and multimodal training data Using Cohen’s Kappa, [14] reports that models trained on larger and more diverse datasets become more human-like in their behaviors. On the other hand, a recent large-scale study using a variation of RSA reported that upgrading from ImageNet-1k to ImageNet21k does not significantly improve alignment to human brain, but object-oriented ImageNet datasets lead to much better alignment than datasets containing only places or faces [11]. Similarly, an ecologically-motivated dataset seems to improve model-brain alignment [23]. Joint vision-language models such as CLIP has also been shown to better predict human brain activity [14, 10, 24].
Similarity measures The reader might have already observed that different similarity measures between representations or behaviors can sometimes yield very different results. Indeed, recent studies warned that different methods could lead to different conclusions [25, 26]. Meanwhile, researches have been examining whether the classic similarity measures are indeed widely applicable: different models result in similar level of linear predicitivity of the brain [18, 11]; RSA is blind to the specific features used to solve a task [27]; later we also show that measures of behavioral similarity may be biased by the choice of decoders [13, 14].
@@ 3 DVC: Quantifying the trial-by-trial consistency of two representations
We develop a new method to evaluate the consistency of two representations. This method is based on a principled generalization of signal detection theory. It enables one to estimate how correlated the decision strategies of two observers are on a classification task. The method is insensitive to the observers’ biases and is not confounded by the behavioral accuracy. It operates at the level of neural representations, and enables one to analyze the internal representation to infer the consistency of the two representations for solving the classification tasks. Thus, the method can quantify task-relevant representational similarity. Compared to methods purely based on behavior [19, 13], it takes advantage of the richness of the internal representations of neural networks and brains. Meanwhile, in contrast to methods for analyzing the similarity of two neural representations (such as representational similarity analysis), our methods focus on the dimensions that are relevant for a behavioral task and is invariant to variability along other task-irrelevant dimensions.
@@ 3.1 Decision variable correlations (DVC) of two neural representations
Signal detection theory is fundamental in the study of perceptual behavior. The idea is that, for binary-choice tasks, observer uses a continuous decision variable (DV) to make a choice (Fig. 1a). Recently, [19] proposed to generalize signal detection theory to study the correlation of decision variables of two observers(Fig. 1b). Their method inferred the decision variable correlation from binary choice data. Here, we develop a simple new strategy to infer the trial-by-trial decision variable correlation from the high-dimensional internal representations (Fig. 1c).
@@
3.2 Implementation of the method
We next discuss practically how we implement the DVC framework to analyze the high-dimensional neural representations from the brains or the deep networks.
Decoding decision variables (DVs) from neural representations For each pair of classes (e.g., cats v.s. dogs), we use Linear Discriminant Analysis (LDA) to find the axis that maximizes class separation to decode the DVs from the brain or model representations. The projection onto the LDA axis reflects the model’s tendency to classify the image as one class versus the other; values near the midpoint indicate greater classification uncertainty. One important practical question is that LDA can be unstable under high dimensions with few samples. The reason is that there are many noisy feature directions with similar class separation, but the projections of image representations along these dimensions can be different. Consequently, even if two models have the same underlying representations, LDA projections may show low correlation. Note that models examined in this paper have a wide range of dimensionality in their penultimate layer ($10^{3}-10^{7}$).
To address this problem, we use dimensionality reduction (e.g., PCA) to reduce the representations to the same number of features before using LDA to decode the underlying DV 11125 PC dimensions. See Appendix C for experiments that demonstrate the robustness of the results.. Importantly, the cross-validated LDA accuracies are high for all representations tested. We measure the similarity between decoded DVs using Pearson Correlation. A DVC value is obtained for each class in each pair of classes. The final reported number is taken as the average of all DVC values.
Normalization for correcting the measurement noise Different systems have different noise levels that limit the correlation that could be reliably decoded. The otherwise perfect correlation between two identical representations would be corrupted by added noise. Low correlation might therefore reflect true underlying dissimilarity or high noise level. To correct for the under-estimation of DVCs due to measurement noise, we develop a split-half procedure to infer the impact of noise.
@@ 3.2 Implementation of the method
This correction removes the bias introduced by independent noise, yielding an unbiased estimate of the true underlying correlation between the latent signals driving $\text{DV}_{A}$ and $\text{DV}_{B}$. 222 The splits are generated randomly. See Appendix A which proves that this recovers the true underlying correlation. See Appendix C for discussions on behaviors of split normalization at boundary conditions.
@@ 4 DVCs reveal the divergence between deep networks and brains
We apply the new DVC-based methodology to examine (i) the trial-by-trial consistency of the neural representation of the macaque high-level visual areas (V4/IT), (ii) the consistency of the neural network models and the IT/V4 neural representations in macaque monkeys, as well as (iii) the consistency of different deep neural network models. We will specifically consider three classes of deep network models: (i) models that were pre-trained on ImageNet-1k using standard network training (i.e., no adversarial training); (ii) "robust models" that were fine-tuned on ImageNet-1k using adversarial training; (iii) "data-rich models" that were pre-trained on even larger datasets such as ImageNet-21k and JFT-300M.
Figure 2: Results on models trained on ImageNet-1k. (a) Heatmap: DVCs inferred for pairs of models. Different colors are used to indicate models from different model families. 15 models are selected to represent this cohort in later analysis. (b) 2D t-SNE embedding of the models using their dissimilarities, measured as $1-DVC$. (c) There is a strong negative correlation between the classification performance (top-1 accuracy) of a network and its DVC correlation to the V4/IT representation. (d) Networks belonging to the same family exhibit higher DVCs compared to those belonging to different model families (p = 1.33e-56).

### v3

@@ Abstract
Previous studies have compared neural activities in the visual cortex to representations in deep neural networks trained on image classification. Interestingly, while some suggest that their representations are highly similar, others argued the opposite. Here, we propose a new approach to characterize the similarity of the decision strategies of two observers (models or brains) using decision variable correlation (DVC). DVC quantifies the image-by-image correlation between the decoded decisions based on the internal neural representations in a classification task. Thus, it can capture task-relevant information rather than general representational alignment. We evaluate DVC using monkey V4/IT recordings and network models trained on image classification tasks. We find that model–model similarity is comparable to monkey-monkey similarity, whereas model–monkey similarity is consistently lower. Strikingly, DVC decreases with increasing network performance on ImageNet-1k. Adversarial training does not improve model–monkey similarity in task-relevant dimensions assessed using DVC, although it markedly increases the model–model similarity. Similarly, pre-training on larger datasets does not improve model–monkey similarity. These results suggest a divergence between the task-relevant representations in monkey V4/IT and those learned by models trained on image classification tasks.
@@ 1 Introduction
Deep learning [bengioRepresentationLearningReview2014, lecunDeepLearning2015] has substantially impacted how neuroscientists construct brain models. For vision neuroscience, deep neural networks offer candidate models for studying the primate ventral pathway [yaminsPerformanceoptimizedHierarchicalModels2014a, khaligh-razaviDeepSupervisedNot2014, yaminsUsingGoaldrivenDeep2016] and, more recently, dorsal pathway [ebrahimpourVentralDorsalNeuralNetworks2020, mineaultYourHeadThere2021, choiDualStreamNeuralNetwork2023]. Early work reported that the representations in convolutional neural networks (CNNs) trained on image categorization tasks can explain a substantial fraction of variance in high-level visual areas, surpassing classic models for these areas [yaminsPerformanceoptimizedHierarchicalModels2014a]. Follow-up research has tested many variants of deep networks on their alignment with the brain using both neural data [muttenthalerHumanAlignmentNeural2023, wangIncorporatingNaturalLanguage2022, conwellWhatCan182023] and behavior data [rajalinghamLargeScaleHighResolutionComparison2018, geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021]. One appealing hypothesis is that deep networks that exhibit higher accuracy and robustness in vision tasks, or trained on larger datasets would better explain visual processing in the brain.
@@
One important question is how to compare deep network models and the brain. One class of methods seeks to quantify the similarity of internal representations between models and brains. This includes methods such as representational similarity analysis (RSA) [kriegeskorteRepresentationalSimilarityAnalysis2008a], linear regression [yaminsUsingGoaldrivenDeep2016, naselarisEncodingDecodingFMRI2011], and generalized shape metrics [williamsGeneralizedShapeMetrics2022a, Harvey2024representational]. More recently, another class of methods that put more emphasis on quantifying the behavioral similarity has been proposed, including Cohen’s Kappa [geirhosAccuracyQuantifyingTrialbytrial2020] and I2n behavioral predictivity [rajalinghamLargeScaleHighResolutionComparison2018, schrimpfBrainScoreWhichArtificial2020]. The goal of these methods is to provide an image-by-image comparison of the decision strategies used by neural networks and the brain. One challenge has been how to properly disentangle the model accuracy, decision biases and decision consistency from behavior [sebastianDecisionvariableCorrelation2018, geirhosAccuracyQuantifyingTrialbytrial2020]. Intriguingly, studies seem to find contradictory trends in model-brain alignment. While some studies suggest that deep learning systems converge to learning a common representation [Huh2024platonic, Shen2025alignment], others suggest that the similarity will not grow indefinitely or has capped [Fel2022harmonizing, Linsley2025better]. The reason for this remains unclear.
@@
Here, we propose a principled approach that combines the merits of model comparisons at the representation and behavior levels. Our method is based on decision variable correlation (DVC) developed to measure the behavioral similarity based on choice data [sebastianDecisionvariableCorrelation2018] in signal detection theory, and we have generalized it to analyze the consistency of high-dimensional neural representations. While prior work developed techniques to estimate DVCs from behavioral data [sebastianDecisionvariableCorrelation2018], we instead infer DVCs from neural representations. Thus, our proposed DVC metric should be interpreted as a measure of the representational similarity. Our approach specifically quantifies the trial-by-trial consistency of two neural representations for solving a classification task, ignoring features that are irrelevant for the task. In doing so, our approach enables one to infer the consistency of the decision strategy of two observers from their internal representations based on the assumption of optimal linear readouts.
Applying our method to compare neural representations for solving image recognition tasks from monkey brains and deep network models led to several interesting findings. In particular, we found that model–model and monkey–monkey similarities are comparable, whereas model–monkey similarity is consistently lower and decreases with increasing ImageNet-1k accuracy. Somewhat surprisingly, this gap is not remedied by adversarial training or training on larger datasets.
@@ 2 Background and relevant work
Community efforts have pushed towards better methods to compare brains and models and for brain-model alignment. Different factors have been hypothesized to be relevant for alignment, including model architecture, robustness, and training data, as summarized below.
@@
Model architecture and scale One hypothesis has been that as models improve in task performance or architectural complexity, their internal representations become more brain-like. The Brain-Scores [schrimpfBrainScoreWhichArtificial2020] of the image classification models were reported to be positively correlated to ImageNet-1k accuracy, although the trend plateaus at higher accuracy. On the other hand, studies using RSA reported that neither model scale nor architecture significantly improved alignment to human behavioral similarity judgments [muttenthalerHumanAlignmentNeural2023]. Another study using RSA reported negative correlation between alignment to human neural activity and model complexity [sartzetakiOneHundredNeural2024]. Studies using Cohen’s Kappa reported that human-model behavioral consistency at the image level remains low despite improved performance on out-of-distribution datasets with scaling[geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021].
@@
Robustness The primate visual system is robust against external and internal noise, prompting the question of whether robustness to adversarial perturbations or corruptions is related to brain-model alignment [dapelloAligningModelMacaque2022]. Recent work proposed that adversarial robustness might promote the learning of representations better aligned with human perception [engstromAdversarialRobustnessPrior2019]. By enforcing alignment with monkey IT representations, models exhibited both enhanced adversarial robustness and increased behavioral alignment with human subjects [dapelloAligningModelMacaque2022]. Another study found that model metamers – artificial stimuli that elicit the same response as natural stimuli, generated by robust models– are more recognizable to humans, but are not themselves predictive of recognizability [featherModelMetamersReveal2023]. However, studies using Cohen’s Kappa report that robust models still diverge from humans in their error patterns[geirhosPartialSuccessClosing2021]. Additionally, there has also been evidence suggesting that even though adversarial training increased model robustness, these robust networks may not use human-like features unless explicitly aligned [Subramanian2023spatial, Linsley2023adversarial].
@@
Rich and multimodal training data Using Cohen’s Kappa, [geirhosPartialSuccessClosing2021] reports that models trained on larger and more diverse datasets become more human-like in their behaviors. On the other hand, a recent large-scale study using a variation of RSA reported that upgrading from ImageNet-1k to ImageNet21k does not significantly improve alignment to human brain, but object-oriented ImageNet datasets lead to much better alignment than datasets containing only places or faces [conwellWhatCan182023]. Similarly, an ecologically-motivated dataset seems to improve model-brain alignment [mehrerEcologicallyMotivatedImage2021]. Joint vision-language models such as CLIP have also been shown to better predict human brain activity [geirhosPartialSuccessClosing2021, wangIncorporatingNaturalLanguage2022, subramaniamRevealingVisionLanguageIntegration2024].
Similarity measures Classic methods such as linear predictivity measure how well neural responses can be predicted from network representation [yaminsPerformanceoptimizedHierarchicalModels2014a]. Representational similarity analysis (RSA) [kriegeskorteRepresentationalSimilarityAnalysis2008a, khaligh-razaviDeepSupervisedNot2014] is a popular approach that quantifies the similarity in the geometrical structure of two representations, and is blind to the specific features used to solve a task [dujmovicInferringDNNBrainAlignment2024]. Other shape metrics such as centered kernel alignment (CKA) and canonical correlation analysis (CCA) also do not specifically address the similarity in task-relevant features. Measures of behavioral similarity such as error consistency are task-focused [geirhosAccuracyQuantifyingTrialbytrial2020], yet they may be overcalibrated on accuracies of the observers and are therefore sensitive to the choice of decoders[geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021] as we will show later. Recent studies highlighted the challenges in the interpretation of results based on these methods [schrimpfBrainScoreWhichArtificial2020, conwellWhatCan182023], e.g., different methods for quantifying the brain-model similarity could lead to different conclusions [soniConclusionsNeuralNetwork2024, sucholutskyGettingAlignedRepresentational2023]. As the field continues to develop methods for analyzing and interpreting model-brain alignments [Barbosa2025quantifying, Chou2025feature], it would be desirable to have principled, task-relevant, accuracy-agnoistic measures to better illustrate the possible divergence between brains and models.
@@ 3 DVC: Quantifying the trial-by-trial consistency of two representations
We develop a new method to evaluate the consistency of two neural representations. This method is based on a principled generalization of signal detection theory. It enables one to estimate how correlated the decision strategies of two observers are for a classification task. The method is robust to the observers’ biases and is not confounded by the behavioral accuracy. It operates at the level of neural representations, and enables one to analyze the internal representation to infer the consistency of the two representations for solving the classification tasks. Thus, the method can quantify task-relevant representational similarity. Compared to methods purely based on behavior [sebastianDecisionvariableCorrelation2018, geirhosAccuracyQuantifyingTrialbytrial2020], it takes advantage of the richness of the internal representations of neural networks and brains. Meanwhile, in contrast to methods for analyzing the similarity of two neural representations (such as representational similarity analysis), our method focuses on the dimensions that are relevant for a behavioral task and is invariant to variability along other task-irrelevant dimensions.
@@ 3.1 Decision variable correlations (DVC) of two neural representations
Signal detection theory is fundamental in the study of perceptual behavior. The idea is that, for binary-choice tasks, observer uses a continuous decision variable (DV) to make a choice (Fig. 1a). Recently, [sebastianDecisionvariableCorrelation2018] proposed to generalize signal detection theory to study the correlation of decision variables of two observers (Fig. 1b). Their method inferred the DVC from binary choice data. Here, we develop a simple new strategy to infer trial-by-trial DVCs from high-dimensional internal representations (Fig. 1c).
@@
3.2 Implementation of the DVC method
We next discuss how we implement the DVC framework to analyze the high-dimensional neural representations from the brains or deep networks.
Decoding decision variables (DVs) from neural representations For each pair of classes (e.g., cats v.s. dogs), we use Linear Discriminant Analysis (LDA) to find the axis that maximizes class separation to decode the DVs from the brain or model representations. The projection onto the LDA axis reflects the model’s tendency to classify the image as one class versus the other; values near the midpoint indicate greater classification uncertainty. One important practical issue is that LDA can be unstable under high dimensions with few samples. The reason is that there are many noisy feature directions with similar class separation, but the projections of image representations along these dimensions can be different. Consequently, even if two models have the same underlying representations, LDA projections may show low correlation. Note that models examined in this paper have a wide range of dimensionality in their penultimate layer ($10^{3}-10^{7}$).
To address this problem, we use dimensionality reduction (e.g., PCA) to reduce the representations to the same number of features before using LDA to decode the underlying DV 11125 PC dimensions. See Appendix C.3 and C.4 for experiments that demonstrate the robustness of the results.. We measure the similarity between decoded DVs using Pearson Correlation. A DVC value is obtained for each class in each pair of classes. The final reported number is taken as the average of all DVC values.
Normalization to address measurement noise in DV  Measurement noise may limit the accuracy of the inferred DVC. The otherwise perfect correlation between two identical representations would be corrupted by adding noise. Low correlation might therefore reflect true underlying dissimilarity or high noise level. To correct for the under-estimation of DVCs due to measurement noise, we develop a split-half procedure to infer the impact of noise.
@@ 3.2 Implementation of the DVC method
This correction removes the bias introduced by independent, additive symmetric noise in the estimated decision variable, yielding an unbiased estimate of the true underlying correlation between the latent signals driving $\text{DV}_{A}$ and $\text{DV}_{B}$  [^2] See Appendix A for proof, and Appendix C for discussions on the validity of DVC under different conditions, including behaviors of split normalization at boundary conditions.. How to best recover DVC under more general noise conditions remains an interesting question for future work.
@@ 4 DVCs reveal the divergence between deep networks and brains
We apply the new DVC-based methodology to examine (i) the trial-by-trial consistency of the neural representation of the high-level visual areas (V4/IT) between macaque brains, (ii) the consistency between individual neural network models & the IT/V4 neural representations, and (iii) the consistency between different deep neural network models. We specifically consider three classes of deep network models: (i) models that were pre-trained on ImageNet-1k using standard network training (i.e., no adversarial training); (ii) "robust models" that were fine-tuned on ImageNet-1k using adversarial training; (iii) "data-rich models" that were pre-trained on even larger datasets such as ImageNet-21k and JFT-300M.
Figure 2: Results on models trained on ImageNet-1k. (a) Heatmap: DVCs inferred for pairs of models. Different colors are used to indicate models from different model families. 15 models are selected to represent this cohort in later analysis. (b) 2D t-SNE embedding of the models using their dissimilarities, measured as 1DVC1-DVC. (c) There is a strong negative correlation between the classification performance (top-1 accuracy) of a network and its DVC correlation to the V4/IT representation. (d) Networks belonging to the same family exhibit higher DVCs compared to those belonging to different model families (p = 1.33e-56).
@@ 4.1 High trial-by-trial consistency of V4 & IT representations across monkey brains
We evaluate the consistency of neural representations in different macaque monkeys. We used the publicly available dataset of objects rendered on naturalistic scenes [majajSimpleLearnedWeighted2015]. In these experiments, they used images from eight classes {animals, boats, cars, chairs, faces, fruits, planes, tables}, with 400 images each, totaling 400 $\times$ 8 = 3200 images. Recordings were taken from V4 and IT areas of two adult macaque monkeys passively viewing these images. The brain representation is taken to be the time-binned spike counts averaged over  50 repeats. 100 neurons from V4 and IT combined were obtained from each monkey.
We first examine the classification accuracy based on V4 and IT response, and find that the accuracy based on LDA is high (0.94 and 0.92, respectively). We combine the neural data from areas V4 and IT, and compute the DVCs. We find that the DVC between the monkeys is about 0.57. We further compute the DVCs for V4 and IT separately, and find the DVC values to be 0.63 and 0.41, respectively. Overall, these results suggest that DVCs across the monkeys’ brain are generally high, implying that the encoding and the decision strategies used by different monkeys are consistent on an image-by-image basis.
We study a set of models (n=43, obtained from Torchvision, [marcelTorchvisionMachinevisionPackage2010]) pretrained on ImageNet-1k, an influential benchmark in image classification. This also offers a fair comparison between models by controlling for confounding factors related to different training data. We use the same neural datasets from [majajSimpleLearnedWeighted2015] as above. We feed the images in [majajSimpleLearnedWeighted2015] into deep vision-based neural networks, subject to standard transforms. The model representation for an image is defined as the activation taken from the penultimate layer – the last layer before the final logit layer.